OpenStudy (anonymous):

The activation energy, Ea. Can be determined from the Arrhenius equation by measuring the rate constants K1 and K2 at 2 different Temperatures T1 and T2 2.303log(k2/k1) = (Ea/R) x ( (T2-T1)/ (T1*T2) ) where T the gas constant is 8.31 mol-1 K-1 Assume that the rate of reaction is ten times faster at 32 than it is at 22. What is the Activation Energy Ea ?

OpenStudy (kittiwitti1):

I think this is a Chemistry problem, not a (specifically) Mathematics problem...

OpenStudy (anonymous):

I feel its a matter of solving for Ea, where T2 = 32 and T1 = 22, and K1 = 10 X k2,

OpenStudy (anonymous):

but I cant rearrange it for Ea

OpenStudy (kittiwitti1):

Does the link help?

OpenStudy (anonymous):

looking at it now, but even then I wouldn't know how to re arrange for Ea lol

OpenStudy (kittiwitti1):

You never know till you try. :3

OpenStudy (unklerhaukus):

\[2.303\log(k_2/k_1) = (E_a/R) \times \frac{T_2-T_1}{T_1\times T_2}\\ 2.303\log(k_2/k_1) \times R= E_a \times \frac{T_2-T_1}{T_1\times T_2}\\ 2.303\log(k_2/k_1) \times R\times\frac{T_1\times T_2}{T_2-T_1}= E_a\]